OpenStudy (anonymous):

Finals this afternoon: Could someone explain this to me? Solve the equation on the interval [-2π,0) 2(sin^2 x)+3sinx+1=0

OpenStudy (anonymous):

let \[\sin x \] be \[a\] \[2a ^{2}+3a+1=0\] \[(2a+1)(a+1)=0\] \[a=-1/2\] and \[a=-1\]so \[\sin x=-1/2\] and \[\sin x=-1\]

OpenStudy (anonymous):

how would i solve sinx=-1/2 and sinx=-1, would i use the unit circle or is there another way to go about solving it?

OpenStudy (anonymous):

you can also use a calculator or special angles for 30-60-90 calculator method \[x=\sin^{-1} (-1/2)=-30=\pi/6\] \[x=\sin^{-1} (-1)=-90=\pi/2\]

OpenStudy (anonymous):

so for your general solution you have \[x=-\pi/6+2\pi.k,\pi \in Z\] and \[x=-\pi/2+2\pi.k \in Z \] \[x=-2\pi,-6/5\pi,-1/2\pi,9/4\]

OpenStudy (anonymous):

okay,,that makes sense, thank you so much for helping me,,,

OpenStudy (anonymous):

my pleasure some steps are omitted in getting x